This is a four-part question.

A. Suppose that the graph of a function f is known.Explain how the graph of
y = 4•f(x) differs from the graph of y = f(4x).

B. Suppose that the graph of a function f is known.Explain how the graph of
y = f(x) - 2 differs from the graph of y = f(x - 2).

C. Suppose that the graph of a function f is known.Explain how the graph of
y = 4•f(x) differs from the graph of y = (1/4)f(x).

D. Suppose that the graph of a function f is known.Explain how the graph of
y = f(x) + 2 differs from the graph of y = f(x + 2).


Answer by ikleyn(52797)

.
This is a four-part question.

A. Suppose that the graph of a function f is known. Explain how the graph of
y = 4•f(x) differs from the graph of y = f(4x).

    Graph y = 4*f(x) is obtained from the parent function graph by stretching  y-axis with coefficient 4.

    Graph y = f(4x)  is obtained from the parent function graph by compressing x-axis with coefficient 4.


B. Suppose that the graph of a function f is known. Explain how the graph of
y = f(x) - 2 differs from the graph of y = f(x - 2).

    Graph y = f(x)-2 is obtained from the parent function graph by translating it 2 units vertically down.

    Graph y = f(x-2)  is obtained from the parent function graph by translating it along x-axis horizontally 2 units right.

----------------------

In the future, do not pack so many questions in one post.

Separate them and submit separately.


////////////////////////////


Comment from student:  It's one question with 4 parts.


My response :   In a joking form,  I could say:  " It is good that the question is only  4  parts,  not  25".


        But if I allow  4-parts question today,
        then tomorrow a  25-parts question will come.


In more serious form,  any human  in  healthy  mind  should be able to separate a multi-parts question assignment
into separate questions/parts.


The rule of this forum is

        One and ONLY ONE problem/question per post.



It is the  RULE,  the  POLICY  and the  REQUIREMENT  of this forum.

It is written in this page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.


It is assumed that you read these rules before posting.

It is also assumed that you do understand what is written in that page and follow the rules.


        This rule is established for benefits of visitors  (to avoid mess in responses).

        Those who violate them,  work against their own interests.